OpenStudy (flowey15):

If the length of the minute hand of the clock in London commonly known as “Big Ben” is 11.25’ from the center of the clock to the tip of the hand, what is the area swept out by the minute hand as it moves 5 minutes of time? A) 10.55 ft B) 28.13 ft C) 33.13 ft D) 66.27 ft

OpenStudy (faiqraees):

Angle created when moved 5 minute = 5/60 *2π Area = 0.5*r^2*theta Area = 0.5*11.25^2^(π/6)